OpenStudy (anonymous):

13 4 7 ÷ 7 is equal to _____. 3 1 3 3 7 8 7

OpenStudy (anonymous):

Anything divided by itself is 1.

OpenStudy (anonymous):

7 divided by 7 is 1

OpenStudy (anonymous):

So what is your answer? @colefoster

OpenStudy (anonymous):

@kaydarro Don't give answers

OpenStudy (anonymous):

why

OpenStudy (vanessaperezxo):

@colefoster 7 divided by 7 equals 1. So, which one do you think is the answer. Test out all the equations to see which one gives the answer.

OpenStudy (anonymous):

i haven't a clue

OpenStudy (vanessaperezxo):

@colefoster pick 2 that you believe could be the answer.

OpenStudy (anonymous):

@kaydarro It's against Open Study's code of conduct, you agreed to it when you joined. http://openstudy.com/code-of-conduct It says: Give Help, Not Answers I will encourage and guide those needing help, and not just give them an answer

OpenStudy (anonymous):

@jaimecontreras Please don't ask your questions on someone else's question.

OpenStudy (anonymous):

oh right thanks

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

ummm a and c

OpenStudy (vanessaperezxo):

Ok, between the 2 which do you feel is the answer?

OpenStudy (amistre64):

if we consider the "\(\div\)" sign as a division operator, then multiplication distributes over division

OpenStudy (anonymous):

a

OpenStudy (amistre64):

\[n\left(\frac{a}{b}\div\frac{c}{d}\right)=\frac{na}{b}\div\frac{nc}{d}\]

OpenStudy (vanessaperezxo):

ok, divide 3 by 1, what does it give you?

OpenStudy (anonymous):

3

OpenStudy (vanessaperezxo):

Ok, so that can't be the answer because they want an equation similar to 7 divided by 7, so that crosses out a

OpenStudy (anonymous):

cross out a

OpenStudy (vanessaperezxo):

Yes, meaning it can't be the answer.

OpenStudy (vanessaperezxo):

Lets do the nest problem. What is 3 divided by 3?

OpenStudy (anonymous):

1

OpenStudy (vanessaperezxo):

Next****

OpenStudy (vanessaperezxo):

Ok, does that give you the answer your looking for?

OpenStudy (vanessaperezxo):

@colefoster

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

so is it b

OpenStudy (vanessaperezxo):

Yes, the answer is B, the reason the answer is be is because 7 divided by 7 = 1 . The equation is asking you to find an equation similar to this one. So, 3 divided by 3 = 1 . Is it similar? Yes, it is. That is why B is the answer.

OpenStudy (anonymous):

thanks

OpenStudy (vanessaperezxo):

Your welcome. Please medal, thanks. :)